I've have found no satisfying answer to this problem.
"Write an equation that is the set of all points in the plane equidistant from the point F(0,3) and the line y=-3"
I've tried and tried to get an answer to this question, going step by step the way it was taught, and cannot arrive at a reasonable answer.
Please help! And Please explain how you arrived at your answer as well, I need to understand this!
Thanks

Answer by dabanfield(803)

The set of points that are equidistant from a line y = -p (called the directrix) and a point (0,p) called the focus is a parabola. The equation can be written as:
x^2 = 4py
In this case p is 3 so the equation of the parabola is:
x^2 = 4*3y
x^2 = 12y
Google "parabola" for more background.